# Random Chowla’s Conjecture for Rademacher Multiplicative Functions

### v1

@@ 5.3. Scales and conditioning
In [MR21] it is shown that the density of $n$ such that $P^{+}(n^{2}+1)>n$ is at least $1/2$ (see also the improvements in [Sah20]). The density of square-free values of $n^{2}+1$ is $\prod_{p\equiv 1\pmod{4}}(1-2/p^{2})\approx 0.894$. It follows that there is a positive proportion of $n$ such that $n^{2}+1$ is square-free and $P^{+}(n^{2}+1)>n$, because the densities add up to more than $1$ and thus the corresponding sets cannot be disjoint. Let the set of such $n\in[N/\log N,N]$ be called $\mathcal{N}$.
@@ 5.5. Creating large fluctuations
We now apply Lemma 5.2 to create large fluctuations for at least one of the first sums $S_{i,1}$. First we calculate the expected variance of $Z_{i}$ over all realizations of $f(q)$ for $q\not\in\mathcal{A}$. $\widetilde{\mathbb{E}}$ denotes conditional expectation, conditioned on the values of $f(q)$ for $q\not\in\mathcal{A}$. We have
(5.11) $\mu_{i}:=\mathbb{E}\mathbb{E}Z_{i}^{2}$ $=\frac{1}{x_{i}}\mathbb{E}\widetilde{\mathbb{E}}S_{i,1}^{2}=\frac{1}{x_{i}}\mathbb{E}\sum_{\begin{subarray}{c}n,m\leq x_{i}\\ p\mid\gcd(n^{2}+1,m^{2}+1)\text{ for some }p\in\mathcal{A}_{i}\\ q\mid(n^{2}+1)/p\implies q\not\in\mathcal{A}\\ q\mid(m^{2}+1)/p\implies q\not\in\mathcal{A}\end{subarray}}f(n^{2}+1)f(m^{2}+1)$
@@
(5.13) $\mathbb{E}(\mathbb{E}Z_{i}^{2})^{2}-\mu_{i}^{2}$ $=\frac{1}{x_{i}^{2}}\mathbb{E}\left(\sum_{\begin{subarray}{c}n,m\leq x_{i}\\ p\mid\gcd(n^{2}+1,m^{2}+1)\text{ for some }p\in\mathcal{A}_{i}\\ q\mid(n^{2}+1)/p\implies q\not\in\mathcal{A}\\ q\mid(m^{2}+1)/p\implies q\not\in\mathcal{A}\end{subarray}}f(n^{2}+1)f(m^{2}+1)\right)^{2}-\mu_{i}^{2}$
(5.14) $=\frac{1}{x_{i}^{2}}\mathbb{E}\sum_{\begin{subarray}{c}n_{1},m_{1},n_{2},m_{2}\leq x_{i}\\ P^{+}(n_{j}^{2}+1)=P^{+}(m_{j}^{2}+1)=:p_{j}\in\mathcal{A}_{i}\,\forall j=1,2\\ q\mid(n_{j}^{2}+1)/p_{j}\implies q\not\in\mathcal{A}\\ q\mid(m_{j}^{2}+1)/p_{j}\implies q\not\in\mathcal{A}\end{subarray}}f(m_{1}^{2}+1)f(n_{1}^{2}+1)f(m_{2}^{2}+1)f(n_{2}^{2}+1)-\mu_{i}^{2}$
(5.15) $=\frac{1}{x_{i}^{2}}\#\left\{m_{j},n_{j}\leq x_{i}:P^{+}(n_{j}^{2}+1)=P^{+}(m_{j}^{2}+1)\,\forall j=1,2,\prod_{j=1}^{2}(m_{j}^{2}+1)(n_{j}^{2}+1)={\raisebox{-1.0pt}{$\square$}}\right\}-\mu_{i}^{2}.$
@@ 6. Acknowledgements
The authors thank Jonathan Bober and Oleksiy Klurman for many helpful conversations throughout the research phase of this project and for their comments on the final draft of this paper. We thank Tim Browning for a fruitful discussion leading to an improvement in Lemma 2.5. The authors also thank Matthew Bisatt, Andrew Granville, Ross Paterson, and Per Salberger for helpful discussions concerning integral points on families of varieties. Much of this work was supported by the Swedish Research Council, grant no. 2021-06594, while the authors were in residence at Institut Mittag-Leffler in Djursholm, Sweden, during the winter semester of 2024; the authors thank the institute and, in particular, the organizers for excellent working conditions. The first author is supported by an NSERC postdoctoral fellowship, application no. PDF – 567986 – 2022; the second author is funded by a University of Bristol PhD scholarship.

### v2

@@ 5.3. Scales and conditioning
In [MR21] it is shown that the density of $n$ such that $P^{+}(n^{2}+1)>n$ is at least $1/2$ (see also the improvements in [Sah20]). The density of $n$ such that $n^{2}+1$ is squarefree is $\prod_{p\equiv 1\pmod{4}}(1-2/p^{2})\approx 0.894$. It follows that there is a positive proportion of $n$ such that $n^{2}+1$ is square-free and $P^{+}(n^{2}+1)>n$, because the densities add up to more than $1$ and thus the corresponding sets cannot be disjoint. Let the set of such $n\in[N/\log N,N]$ be called $\mathcal{N}$.
@@ 5.5. Creating large fluctuations
We now apply Lemma 5.2 to create large fluctuations for at least one of the first sums $S_{i,1}$. First, we calculate the expected variance of $Z_{i}$, which is equal to the expectation of $\mathbb{E}Z_{i}^{2}$ over all realizations of $f(q)$ for $q\not\in\mathcal{A}$. We denote conditional expectation by $\widetilde{\mathbb{E}}$, conditioned on the values of $f(q)$ for $q\not\in\mathcal{A}$, and recall that $\mathbb{E}Z_{i}^{2}$ is equal to the (appropriately normalized) conditional variance of $S_{i,1}$. We have
(5.11) $\beta_{i}:=\mathbb{E}\mathbb{E}Z_{i}^{2}$ $=\frac{1}{x_{i}}\mathbb{E}\widetilde{\mathbb{E}}S_{i,1}^{2}=\frac{1}{x_{i}}\mathbb{E}\sum_{\begin{subarray}{c}n,m\leq x_{i}\\ p\mid\gcd(n^{2}+1,m^{2}+1)\text{ for some }p\in\mathcal{A}_{i}\\ q\mid(n^{2}+1)/p\implies q\not\in\mathcal{A}\\ q\mid(m^{2}+1)/p\implies q\not\in\mathcal{A}\end{subarray}}f(n^{2}+1)f(m^{2}+1)$
@@
(5.13) $\mathbb{E}(\mathbb{E}Z_{i}^{2})^{2}-\beta_{i}^{2}$ $=\frac{1}{x_{i}^{2}}\mathbb{E}\left(\sum_{\begin{subarray}{c}n,m\leq x_{i}\\ p\mid\gcd(n^{2}+1,m^{2}+1)\text{ for some }p\in\mathcal{A}_{i}\\ q\mid(n^{2}+1)/p\implies q\not\in\mathcal{A}\\ q\mid(m^{2}+1)/p\implies q\not\in\mathcal{A}\end{subarray}}f(n^{2}+1)f(m^{2}+1)\right)^{2}-\beta_{i}^{2}$
(5.14) $=\frac{1}{x_{i}^{2}}\mathbb{E}\sum_{\begin{subarray}{c}n_{1},m_{1},n_{2},m_{2}\leq x_{i}\\ P^{+}(n_{j}^{2}+1)=P^{+}(m_{j}^{2}+1)=:p_{j}\in\mathcal{A}_{i}\,\forall j=1,2\\ q\mid(n_{j}^{2}+1)/p_{j}\implies q\not\in\mathcal{A}\\ q\mid(m_{j}^{2}+1)/p_{j}\implies q\not\in\mathcal{A}\end{subarray}}f(m_{1}^{2}+1)f(n_{1}^{2}+1)f(m_{2}^{2}+1)f(n_{2}^{2}+1)-\beta_{i}^{2}$
(5.15) $=\frac{1}{x_{i}^{2}}\#\left\{m_{j},n_{j}\leq x_{i}:P^{+}(n_{j}^{2}+1)=P^{+}(m_{j}^{2}+1)\,\forall j=1,2,\prod_{j=1}^{2}(m_{j}^{2}+1)(n_{j}^{2}+1)={\raisebox{-1.0pt}{$\square$}}\right\}-\beta_{i}^{2}.$
@@ 6. Acknowledgements
The authors thank: Jonathan Bober and Oleksiy Klurman for many helpful conversations throughout the research phase of this project and for their comments on the final draft of this paper; Tim Browning for a fruitful discussion leading to an improvement in Lemma 2.5; Matthew Bisatt, Andrew Granville, Ross Paterson, and Per Salberger for helpful discussions concerning integral points on families of varieties; and the anonymous referee for their diligence in reviewing this manuscript. Much of this work was supported by the Swedish Research Council, grant no. 2021-06594, while the authors were in residence at Institut Mittag-Leffler in Djursholm, Sweden, during the winter semester of 2024; the authors thank the institute and, in particular, the organizers for excellent working conditions. The first author is supported by an NSERC postdoctoral fellowship, application no. PDF – 567986 – 2022; the second author is funded by a University of Bristol PhD scholarship.